The perimeter of a triangle is 450 m and it's sides are in the ratio 6:7:8. Find the area of the triangle.
Answer by dkppathak(439)

The perimeter of a triangle is 450 m and it's sides are in the ratio 6:7:8. Find the area of the triangle.
let Sides of triangle are 6x,7x,8x and perimetre is 450
6x+7X+8X=450
s=(a+b+C)/2
21x=450
x=450/21
=150/7 than side are 900/7,1050/7 ,and 1200/7
s= 3150/7x2=1575/7=225
by using heron formula for area of scalene triangle
sqrt[s(s-a)(s-b)(s-c)
sqrt(225x675x525x375/7x7x7)
=9336.6562
answer